# Baryon-number-violating nucleon decays in sterile neutrino effective field theories

### v1

@@ I Introduction
Many theoretical frameworks predict baryon-number-violating (BNV) nucleon decays, including Grand Unified Theories (GUTs) Georgi:1974sy ; Fritzsch:1974nn ; Langacker:1980js ; Nath:2006ut and supersymmetry (SUSY) Farrar:1978xj ; Dimopoulos:1981zb ; Sakai:1981gr . BNV nucleon decays are also motivated from the viewpoint of effective field theories (EFTs) deGouvea:2014lva ; Heeck:2019kgr ; Girmohanta:2019xya ; Antusch:2020ztu ; He:2021sbl ; Fajfer:2023gfi ; Beneito:2023xbk ; Gargalionis:2024nij ; Beneke:2024hox (see Refs. FileviezPerez:2022ypk ; Dev:2022jbf ; Ohlsson:2023ddi for recent reviews), since the baryon number is an accidental symmetry of the SM. The large lower bounds on the lifetime of the proton place stringent limits on the very massive mediators and the NP energy scale of the BNV EFT operators. The lower limits on the NP energy scale are found to be $\sim 10^{15}$ GeV and $10^{11}$ GeV for dimension-6 or dimension-7 BNV operators in the SM effective field theory (SMEFT) Beneito:2023xbk , respectively.
@@
Table 1: Baryon-number-violating two-body nucleon decays with a neutrino $\nu$ or antineutrino $\bar{\nu}$ and a pseudoscalar meson $P$ in the final states. As the nature of the final state neutrino is not experimentally determined, we collectively denote them by $\nu$. All limits are at 90% C.L. .
This paper is organized as follows. In Sec. II, we describe the effective field theory frameworks for baryon number violation at the relevant energy scales. The formulas of BNV nucleon decay rates are presented in Sec. III. In Sec. IV we discuss the nucleon decay searches with missing energy and show the numerical results for the two-body BNV nucleon decays to neutrino. Our main conclusions are summarized in Sec. V.
II Sterile neutrinos and BNV operators in effective field theories
@@
There is an active experimental program dedicated to the search for sterile neutrinos or heavy neutral leptons, as recently summarized in Ref. Abdullahi:2022jlv . Demanding the sterile neutrinos to contribute to active neutrino masses via the seesaw mechanism Minkowski:1977sc translates into a lower bound on the mass of sterile neutrinos. Big bang nucleosynthesis (BBN) imposes a lower bound on the lifetime of sterile neutrinos of $\tau_{N}<0.02s$ Boyarsky:2020dzc which translates into a lower bound on the active-sterile mixing angle. Beam dump and collider experiments provide an upper limit on the active-sterile mixing angle. The BBN constraint together with the kinematic threshold of $K\to\pi\nu N$ excludes sterile neutrinos lighter than 330–360 MeV except for a small mass window between 120–140 MeV. Bondarenko:2021cpc
II.1 BNV operators in SMEFT, $\nu$SMEFT and LEFT
@@
where $C_{i}$ denotes the WCs for the BNV operators and $p,r,s,t$ refer to the generation indices of the fermions to the order in the operators of Tab. 2. We work in the up-type quark basis and include the dimension-6 SMEFT operators with $|\Delta(B-L)|=0$ (top-left) Jenkins:2017jig , dimension-7 operators with $|\Delta(B-L)|=2$ (bottom-left) Liao:2020zyx , dimension-6 $\nu$SMEFT operators with $|\Delta(B-L)|=0$ (top-right) Li:2020lba ; Li:2021tsq and dimension-7 $\nu$SMEFT operators with $|\Delta(B-L)|=2$ (bottom-right) Liao:2016qyd ; Li:2020lba ; Li:2021tsq . The operators with checkmark $\checkmark$” are relevant for the nucleon decay to neutrino and we will focus on them below.
@@
II.2 Matching of SMEFT and $\nu$SMEFT to LEFT
@@ II.2.2 $\nu$SMEFT operators
II.3 Renormalization group evolution
@@
We start from the RGEs of the SMEFT/$\nu$SMEFT WCs which are used to run the operators from UV scale $\Lambda$ down to the EW scale $m_{Z}$. The approximate RGEs of the dimension-6 SMEFT WCs are Alonso:2014zka
@@
where $q_{u}$ ($q_{d}$) denotes the electric charge of up-type (down-type) quark. We use them to evolve the WCs down to the nucleon scale $\mu=2$ GeV.
II.4 Matching to the BChPT
@@ II.5 Discussion on UV models
III Baryon-number-violating Lagrangian and nucleon decay rates
III.1 Effective Lagrangian and nucleon decay rate
@@
III.2 Description of our calculation method
@@
IV Numerical results
IV.1 Overview of nucleon decay searches with missing energy
@@
IV.2 Numerical results for the two-body BNV nucleon decays
@@ IV.2.2 $\nu$SMEFT only
Figure 5: Bounds on three illustrative WCs and UV energy scales in $\nu$SMEFT with $s=4$ derived from different processes as functions of the sterile neutrino mass $m_{N}$. See the discussion in Sec. IV.1 for caveats in reinterpreting the experimental limits. In particular, the $p\to\pi^{+}\nu$ limit actually doesn’t apply for sterile neutrino mass above 665 MeV.
IV.2.3 Both SMEFT and $\nu$SMEFT
@@ Appendix A Electroweak contributions to dimension-7 $\nu$SMEFT operators
$\dot{C}=\frac{\mu}{16\pi^{2}}\frac{dC}{d\mu}=\delta C_{,1}-\frac{1}{2}\sum_{i\in I|J}\delta Z_{\phi_{i},1}C+\dots\;,$ (101)

### v2

@@ 1 Introduction
Many theoretical frameworks predict baryon-number-violating (BNV) nucleon decays, including Grand Unified Theories (GUTs) Georgi:1974sy ; Fritzsch:1974nn ; Langacker:1980js ; Nath:2006ut and supersymmetry (SUSY) Farrar:1978xj ; Dimopoulos:1981zb ; Sakai:1981gr . BNV nucleon decays are also motivated from the viewpoint of effective field theories (EFTs) deGouvea:2014lva ; Heeck:2019kgr ; Girmohanta:2019xya ; Antusch:2020ztu ; He:2021sbl ; Fajfer:2023gfi ; Beneito:2023xbk ; Gargalionis:2024nij ; Beneke:2024hox ; Gisbert:2024sjw (see Refs. FileviezPerez:2022ypk ; Dev:2022jbf ; Ohlsson:2023ddi for recent reviews), since the baryon number is an accidental symmetry of the SM. The large lower bounds on the lifetime of the proton place stringent limits on the very massive mediators and the NP energy scale of the BNV EFT operators. The lower limits on the NP energy scale are found to be $\sim 10^{15}$ GeV and $10^{11}$ GeV for dimension-6 or dimension-7 BNV operators in the SM effective field theory (SMEFT) Beneito:2023xbk , respectively.
@@
Table 1: Lower limits on the lifetimes of baryon-number-violating two-body nucleon decays with a neutrino $\nu$ or antineutrino $\bar{\nu}$ and a pseudoscalar meson $P$ in the final states. As the nature of the final state neutrino is not experimentally determined, we collectively denote them by $\nu$. All limits are at 90% C.L. .
This paper is organized as follows. In Sec. 2, we describe the effective field theory frameworks for baryon number violation at the relevant energy scales. The formulas of BNV nucleon decay rates are presented in Sec. 3. In Sec. 4 we discuss the nucleon decay searches with missing energy and show the numerical results for the two-body BNV nucleon decays to neutrino. Our main conclusions are summarized in Sec. 5.
2 Sterile neutrinos and BNV operators in effective field theories
@@
There is an active experimental program dedicated to the search for sterile neutrinos or heavy neutral leptons, as recently summarized in Ref. Abdullahi:2022jlv . Demanding the sterile neutrinos to contribute to active neutrino masses via the seesaw mechanism Minkowski:1977sc translates into a lower bound on the mass of sterile neutrinos. Big bang nucleosynthesis (BBN) imposes an upper bound on the lifetime of sterile neutrinos of $\tau_{N}<0.02s$ Boyarsky:2020dzc which translates into a lower bound on the active-sterile mixing angle. Beam dump and collider experiments provide an upper limit on the active-sterile mixing angle. The BBN constraint together with the kinematic threshold of $K\to\pi\nu N$ excludes sterile neutrinos lighter than 330–360 MeV except for a small mass window between 120–140 MeV. Bondarenko:2021cpc
2.1 BNV operators in SMEFT, $\nu$SMEFT and LEFT
@@
where $C_{i}$ denotes the WCs for the BNV operators and $p,r,s,t$ refer to the generation indices of the fermions to the order in the operators of Tab. 2. We work in the up-type quark basis  [^2] In the up-type quark basis, the up-type quark mass matrix is assumed to be diagonal and the CKM matrix diagonalizes the down-type quark mass matrix. and include the dimension-6 SMEFT operators with $|\Delta(B-L)|=0$ (top-left) Jenkins:2017jig , dimension-7 operators with $|\Delta(B-L)|=2$ (bottom-left) Liao:2020zyx , dimension-6 $\nu$SMEFT operators with $|\Delta(B-L)|=0$ (top-right) Li:2020lba ; Li:2021tsq and dimension-7 $\nu$SMEFT operators with $|\Delta(B-L)|=2$ (bottom-right) Liao:2016qyd ; Li:2020lba ; Li:2021tsq . The operators with checkmark $\checkmark$” are relevant for the nucleon decay to neutrino and we will focus on them below.
@@
For illustration, the operator $QQLQ$ has an index symmetry described by the sum of the 3 Young tableaux. It allows $1+8+10=19$ possible index combinations, where we used the same order as the Young tableaux. Hence, there have to be $83^{3}-19=$ 8 vanishing operators.
2.2 Matching of SMEFT and $\nu$SMEFT to LEFT
@@ 2.2.2 $\nu$SMEFT operators
2.3 Renormalization group evolution
@@
We start from the RGEs of the SMEFT/$\nu$SMEFT WCs which are used to run the operators from UV scale $\Lambda$ down to the EW scale $m_{Z}$. The approximate RGEs of the dimension-6 SMEFT WCs are  [^3] We ignore Yukawa terms of lepton, light quarks and sterile neutrino, and only keep gauge contributions and top Yukawa terms in RGEs. Alonso:2014zka
@@
where $q_{u}$ ($q_{d}$) denotes the electric charge of up-type (down-type) quark. We use them to evolve the WCs down to the nucleon scale $\mu=2$ GeV which is used in lattice calculation for hadronic parameters Yoo:2021gql .
2.4 Matching to the BChPT
@@
Here and above, $u_{pL}$ and $u_{BL}$ ($u_{BR}$) stand for the baryon states projected by chirality projector $P_{L}$ ($P_{R}$).
@@ 2.5 Discussion on UV models
3 Baryon-number-violating Lagrangian and nucleon decay rates
3.1 Effective Lagrangian and nucleon decay rate
@@
3.2 Description of our calculation method
@@
4 Numerical results
4.1 Overview of nucleon decay searches with missing energy
@@
4.2 Numerical results for the two-body BNV nucleon decays
@@ 4.2.2 $\nu$SMEFT only
Note that the couplings and decay rates for BNV decays are both given with respect to the LEFT WCs, as shown in Tabs. 4 and 5. However, Tab. 2 shows that SMEFT and $\nu$SMEFT operators with the same $B-L$ number have opposite neutrino chiralities. The LEFT operators are all expressed by $\nu_{L}$ and its charge conjugate state $\nu_{L}^{c}$. Thus, as shown in Tab. 3, dimension-6 (7) SMEFT operators with $\Delta(B-L)=0$ ($|\Delta(B-L)|=2$) and dimension-7 (6) $\nu$SMEFT operators with $|\Delta(B-L)|=2$ ($\Delta(B-L)=0$) are matched to the same LEFT operators but with different neutrino flavor indices. Note that $\mathcal{O}_{ddu}^{S,RL}$ can only be induced by $\mathcal{O}_{HdNQ}$ operator. One can see that $\Delta(B-L)=0$ decays $N\to M\overline{\nu}$ ($|\Delta(B-L)|=2$ decays $N\to M\nu$) are indeed induced by LEFT operators with $\Delta(B-L)=0$ ($|\Delta(B-L)|=2$).
@@
Figure 5: Bounds on three illustrative WCs and UV energy scales in $\nu$SMEFT with $s=4$ derived from different processes as functions of the sterile neutrino mass $m_{N}$. See the discussion in Sec. 4.1 for caveats in reinterpreting the experimental limits. In particular, the $p\to\pi^{+}\nu$ limit actually doesn’t apply for sterile neutrino mass above 665 MeV.
4.2.3 Both SMEFT and $\nu$SMEFT
@@ Appendix A Electroweak contributions to dimension-7 $\nu$SMEFT operators
$\dot{C}=16\pi^{2}\mu\frac{dC}{d\mu}=\delta C_{,1}-\frac{1}{2}\sum_{i\in I|J}\delta Z_{\phi_{i},1}C+\dots\;,$ (101)